OpenStudy (anonymous):

The area of a rectangle is 48 square centimeters. The length of the rectangle is 8 centimeters longer than the width. Which equation models this situation? A) w + 8 = 48 B) w(w + 8) = 48 C) 8w = 48 D) w + 8w = 48

jhonyy9 (jhonyy9):

the length is 8 cm longer than the width so this mean l=w+8

OpenStudy (anonymous):

B.

OpenStudy (campbell_st):

its not B

OpenStudy (anonymous):

i still don't get it

OpenStudy (anonymous):

it is B..... the formula to get an area of a rectangle is: length (L) x Width (W) = Area (A) LxW=A We know that A=48 and we also know that the Length(L) is 8cm longer than the Width(W) which means: L= 8+W Now we have two equations: 1st eqaution: 48=LxW 2nd equation: L=8+W We know what the Length equals by looking at the second equation so use that and plug what L equals into the first equation.So now we would have: 48=(8+W)W Which BY THE WAY is B. @campbell_st @NormaV

OpenStudy (anonymous):

jhonyy9, i don't understand what you mean

OpenStudy (anonymous):

are you ignoring my post.... i literally explained how you solve this whole equation....

OpenStudy (anonymous):

If you read what i wrote then you'll understand, I'm out of here. you're welcome by the way. @NormaV

OpenStudy (anonymous):

no i'm sorry my computer is lagging